a car is travelling with a speed of 90km/hr reduces its speed 36km/hr.while travelling a distance of speed 100 m. find the deceleration produced?
Answer by Alan3354(69443)

a car is travelling with a speed of 90km/hr reduces its speed 36km/hr.while travelling a distance of speed [sic] 100 m. find the deceleration produced?
--------------
a car is travelling with a speed of 90km/hr reduces its speed 36km/hr.while travelling a distance of 100 m. find the deceleration produced?
--------------
90 km/hr = 25 m/sec
36 km/hr = 10 m/sec
-----
The avg speed over the 100 meters = (25 + 10)/2 = 17.5 m/sec
t = 100/17.5 = 40/7 seconds
-----
a = (10 - 25)/(40/7) = -21/8 m/sec/sec
a = -2.625 m/sec/sec